the sum of the ages of the father and his son is 40 years. if the they both live on till the son becomes as old as the father is now, the sum of their age will be 96 years. find their present ages.
Answer by macston(5194)

.
F=father's age; S=son's age
F+S=40
F=40-S
.
x=number years till son is as old as father
(F+x)+(S+x)=96
(F+S)+2x=96
40+2x=96
2x=56
x=28
F=S+28
40-S=S+28
12=2S
6=S ANSWER 1: The son is currently 6 years old.
F=40-S=40-6=34 ANSWER 2: The father is currently 34 years old.